OpenStudy (anonymous):

Find the sum and choose the correct answer. 2n 3 + 4n 2 - 7 and -n 3 + 8n - 9 -n3 - 4n2 - 8n - 16 n3 + 4n2 + 8n - 16 n3 + 4n2 + 8n + 16 3n3 + 4n2 + 8n - 2

OpenStudy (anonymous):

@Abhisar

geerky42 (geerky42):

what do you think?

geerky42 (geerky42):

just combine like terms

OpenStudy (anonymous):

A?

geerky42 (geerky42):

let's do the sum: \[\Large (2n^3 + 4n 2 - 7) + (-n^3 + 8n - 9)\]

geerky42 (geerky42):

We can see that \(2n^3\) and \(-n^3\) both have like term ( \(n^3\) ) so what's \(2n^3+(-n^3)\)?

OpenStudy (anonymous):

im so confused gosh i hate math

OpenStudy (texaschic101):

try this... like terms are gonna have the same letter and exponent 2n^3 - n^3 = 4n^2 -- no other like term 8n -- no other like term -9 - 7 = so what is 2n^3 - n^3......-n^3 is the same as -1n^3 if that helps. and what is -9 - 7 ? can you answer these 2 questions ?

OpenStudy (anonymous):

no i dont like algebra im passing by a 60%

OpenStudy (texaschic101):

2n^3 - 1n^3.....the n^3 stays the same, so what is 2 - 1 ?

OpenStudy (anonymous):

2-1=1

OpenStudy (texaschic101):

let me just show you... 2n^3 - n^3 = n^3 4n^2 8n -9 - 7 = -16 now put it all together.. n^3 + 4n^2 + 8n - 16 please ask if you have any questions

OpenStudy (texaschic101):

2n^3 - n^3 is the same as 2n^3 - 1n^3 = 1n^3...which is the same as n^3